# Blending Complementary Memory Systems in Hybrid Quadratic-Linear Transformers

### v1

@@ Abstract
We develop hybrid memory architectures for general-purpose sequence processing neural networks, that combine key-value memory using softmax attention (KV-memory) with dynamic synaptic memory through fast-weight programming (FW-memory)—the core principles of quadratic and linear transformers, respectively. These two memory systems have complementary but individually limited properties: KV-memory offers precise retrieval but is constrained by quadratic complexity in sequence length, while FW-memory supports arbitrarily long sequences and enables more expressive computation but sacrifices precise recall. We propose and compare three methods to blend these two systems into a single memory system to leverage the strengths of both. We conduct experiments on general language modeling and retrieval tasks by training 340M- and 1.3B-parameter models from scratch, as well as on synthetic algorithmic tasks designed to precisely illustrate the benefits of certain hybrid methods over others. We also evaluate our hybrid memory systems on reinforcement learning in partially observable environments. Overall, we demonstrate how a well-designed hybrid can overcome the limitations of its individual components, offering new insights into the design principle of neural memory systems. [^1] Our code is public: https://github.com/kazuki-irie/hybrid-memory.
@@ 1 Introduction
Modern transformers come in two flavors. The standard or “quadratic transformer” (QT) [1] leverages softmax attention over explicit key-value memory storage which enables precise memory retrieval [2], while the length of sequences they can handle is limited in practice due to its quadratic computational complexities (as reflected in our model terminology). On the other hand, “linear transformers” (LT), also known as fast weight programmers (FWPs) [3, 4] sacrifice the precision of the softmax function (the complexity bottleneck of QT), but in return they support the processing of arbitrarily long sequences. Furthermore, some of their variants (called DeltaNet models [5, 6, 7]) enable more expressive computations (e.g., certain types of state-tracking), which QTs cannot perform [8, 7, 9].
@@
We propose and compare three blending methods that are conceptually well-motivated, and empirically evaluate them to ultimately converge on a single architecture. We conduct experiments to test general language modeling and in-context retrieval abilities (using the standard lm-evaluation-harness [12]), by training 340M- and 1.3B-parameter language models from scratch using 15B tokens of the HuggingFace FineWeb-Edu dataset [13]. We also conduct experiments on synthetic algorithmic tasks (parity and modular arithmetics) [8] to determine the type of hybrid methods that can preserve the expressivity advantages of DeltaNet (which we use as the LT/FWP component of our hybrid)—a critical discussion and evaluation that are completely absent in prior discussions on quadratic and linear transformer hybrids [14, 15]. Overall, we demonstrate the benefits of hybrid model designs that can take the full advantage of the latest LT/FWP models, and show how a well-designed hybrid can mitigate the limitations of its individual components; this offers novel insights into the design principle of neural memory systems.
Finally, all our models are scalable as they are compatible with the algorithms and implementations for highly efficient quadratic and linear transformers that have been proposed in prior work on flash-attention [16] and flash-linear-attention [17]. We will make all our experimental settings and training recipes available on a public GitHub repository upon acceptance (e.g., the language modeling part is based on flame [18]).
@@ 2.1 Key-value memory & Softmax attention
where ${\bm{q}}_{t},{\bm{k}}_{t}\in\mathbb{R}^{d_{\text{key}}}$, ${\bm{v}}_{t}\in\mathbb{R}^{d_{\text{out}}}$ (with [][][ ] denoting vector concatenation), ${\bm{W}}^{\text{slow}}\in\mathbb{R}^{(2d_{\text{in}}+d_{\text{out}})\times d_{\text{in}}}$ is the trainable weight matrix (meaning of the superscript “slow” becomes clear in Sec. 2.2), $\oplus$ as in ${\bm{K}}_{t-1}\oplus{\bm{k}}_{t}$ denotes concatenation of vector ${\bm{k}}_{t}\in\mathbb{R}^{d_{\text{key}}}$ to matrix ${\bm{K}}_{t-1}\in\mathbb{R}^{d_{\text{key}}\times(t-1)}$ along the time dimension, yielding ${\bm{K}}_{t}\in\mathbb{R}^{d_{\text{key}}\times t}$ (${\bm{K}}_{0}$ and ${\bm{V}}_{0}$ are initially empty), and $\mathrm{softmax}$ is applied along the time dimension. We omit the $1/\sqrt{d_{\text{key}}}$ scaling inside $\mathrm{softmax}$, as well as the output projection, which are irrelevant to our discussion.
Eq. 3 is the so-called softmax attention operation; the query vector from the current step $t$ is compared to the keys from all the time steps through dot product; these similarity scores are sharpened through the softmax function (crucial for precise retrieval), and the resulting scores are used as coefficients to compute the weighted average of all the value vectors to produce the output. As can be seen in Eqs. 2, the sizes of key and value memory storage linearly grow with the time step (i.e., sequence length), resulting in quadratic complexity w.r.t. sequence length in the attention computation in Eq. 3. Consequently, practical self-attention requires to predetermine a maximum context length, i.e., size of the sliding window, and any input that falls outside the window is discarded. On the other hand, training of such a model can be highly efficient on modern hardware as the computation above is parallelizable over time steps/sequence elements.
@@ 2.2 Fast weight memory & Linear attention and delta rule
where the dimension of ${\bm{W}}^{\text{slow}}$ is increased to be $\in\mathbb{R}^{(2*d_{\text{key}}+d_{\text{out}}+1)\times d_{\text{in}}}$, i.e., one extra output dimension ($+1$) is introduced to produce an addition variable $\beta_{t}\in\mathbb{R}$. $\sigma$ is an activation function applied to $\beta_{t}$, which is typically sigmoid or $2$ times sigmoid (the factor $2$ introduces negative eigenvalues in the state transition matrix enabling strong state-tracking abilities [8]). The choice of $\phi$ is particularly important for stability when the delta rule is used [5]; here we define $\phi$ as element-wise sigmoid linear unit (SiLU; $x$ times sigmoid) followed by $L_{2}$ normalization as in Yang et al. [23]. Eq. 8 corresponds to a rank-one update of the fast weight matrix, from ${\bm{W}}_{t-1}$ to ${\bm{W}}_{t}$, through the delta learning rule [28], where the slow net-generated variables, ${\bm{v}}_{t}$, $\phi({\bm{k}}_{t})$, and $\sigma(\beta_{t})$, play the role of target, input, and learning rate of the delta rule, respectively.
@@ 3.1 Delayed-Streaming HQLT
Let $\ominus$ denote ‘remove’ operation taking a matrix and one of its column vectors to be removed as arguments. $S$ denotes the window size, i.e., the number of key-value pairs KV-memory can store.
At every time step $t$, an HQLT takes an input ${\bm{x}}_{t}\in\mathbb{R}^{d_{\text{in}}}$, and produces an output ${\bm{y}}_{t}\in\mathbb{R}^{d_{\text{out}}}$, while maintaining two types of memory: (1) KV-memory, denoted as two matrices ${\bm{K}}_{t-1}\in\mathbb{R}^{d_{\text{key}}\times S}$ and ${\bm{V}}_{t-1}\in\mathbb{R}^{d_{\text{out}}\times S}$, and (2) FW-memory, denoted as a matrix ${\bm{W}}_{t-1}\in\mathbb{R}^{d_{\text{out}}\times d_{\text{in}}}$. As in DeltaNet (Sec. 2.2), the trainable parameter matrix is ${\bm{W}}^{\text{slow}}\in\mathbb{R}^{(2*d_{\text{key}}+d_{\text{out}}+1)\times d_{\text{in}}}$, and variables are first generated from the input as follows:
@@ 3.2 Delayed-Chunk HQLT
While Delayed-Streaming HQLT above seems natural from the perspective of streaming causal sequence processing, another “delayed” variation can be motivated by the chunk-wise training algorithm of Eq. 6 (left). In fact, an HQTL can be naturally obtained by introducing the softmax activation in the intra-chunk attention operation of Eq. 6 (left; the second term), i.e., $(\mathbf{K}_{\mathbf{n}}^{\top}\mathbf{Q}_{\mathbf{n}}\odot\mathbf{M})$; and by using this chunk-wise processing for inference too. This approach is naturally compatible with efficient training: we can apply flash-linear-attention for inter-chunk attention computation, and flash-attention to the intra-chunk attention with softmax. We call this approach “Delayed-Chunk HQLT”.
@@ 4.1 Evaluating general language modeling capabilities
We begin with evaluating general language modeling capabilities of the proposed HQTL models as a general sanity check. We train language models with either 340M or 1.3B trainable parameters from scratch on 15B tokens of the HuggingFace FineWeb-Edu dataset [13]. These choices mostly follow the configurations of the recent work on LTs/DeltaNet [22, 23, 8] and their most recent recommendations available on the flame repository [17], including the choice of tokenizer (fla-hub/transformer-1.3B-100B). The subword-unit vocabulary size is 32K for all models. All models have 24 layers. The baseline Transformer architecture is from [32] (denoted as Transformer++, following prior convention) and the DeltaNet configuration is from [23]; both models/configurations are used in prior work [23, 8]. Note, however, that for fair comparison with the baseline quadratic transformer, we do not use short-window convolution in the DeltaNet, unlike in recent work [23].
@@
    window ×$\times 2$ = 128 26.6 27.8 35.6 68.1 41.7 51.2 61.4 30.1 48.0
    window ×$\times 4$ = 256 26.7 27.7 35.4 67.0 42.2 52.7 59.8 29.0 47.7
    window ×$\times 8$ = 512 27.0 28.0 35.9 66.3 41.3 53.2 60.1 29.0 47.6
@@ 4.2 Evaluating Expressivity
Here we compare HQLT variants’ abilities to leverage expressivity advantages of the FW-memory component, DeltaNet. We evaluate models on two popular regular language recognition tasks, parity and modular arithmetic without brackets, which are well-known tasks that quadratic transformers and other linear recurrent models fail (due to lack of expressivity), while more expressive, DeltaNet performs well [8]. We mainly follow Grazzi et al. [8]’s experimental settings: we use 2-layer and 3-layer models for parity and modular arithmetic, respectively; training and test sequences have lengths from 3 up to 40, and from 40 to 256, respectively. The window size for all HQLTs is set to 16. We refer to Appendix A.2 for further experimental details.
Table 3 shows the results. As expected, the two “delayed” HQLT models, Delayed-Stream and Delayed-Chunk, fail at these tasks just like the quadratic transformer. In these “delayed” variations, FW-memory only gets access to an input after a delay of a few time steps corresponding to the quadratic attention window size; this could be too late to leverage FW-memory’s capabilities for these tasks which require state-tracking in “real time”. In contrast, HQLT Synchronous, which has no such limitation, successfully solves these tasks.
@@
Given that the delayed versions have no other strong advantage over the synchronous version (apart from the conceptual elegance in introducing explicit division of labors over time), we consider this result as a strong argument to prefer the Synchronous variant over other blending strategies for HQLT. This also allows us to conclude that the design of “Infini-attention” from prior work [15], which makes use of the Delayed-Chunk strategy, is a sub-optimal choice in light of expressivity.
Table 4: Evaluating Hybrid Quadratic-Linear Transformers (HQLTs) on retrieval-intensive tasks. Both the 340M/1.3B models are trained for 15B tokens. The default window size for KV-memory is 64 tokens in HQLTs. Transformer++’s context size is 2048.
SWDE SQuAD FDA Avg.
acc $\uparrow$ acc $\uparrow$ acc $\uparrow$
@@
   sum mixer 13.3 26.2 12.6 17.4
   dynamic scalar (0.4M) 21.1 27.7 11.4 20.1
   dynamic vector (25M) 20.0 28.4 10.9 19.8
    window ×$\times 2$ = 128 16.9 30.5 17.9 21.8
    window ×$\times 4$ = 256 18.6 35.6 15.1 23.1
    window ×$\times 8$ = 512 22.9 35.7 17.3 25.3
    window ×$\times 16$ = 1024 34.0 37.1 50.1 40.4
@@
Transformer++ 53.7 41.5 64.7 53.3
DeltaNet 32.9 29.9 23.6 28.8
HQLT Synchronous 31.9 31.2 30.2 31.1
@@ 4.3 Evaluating in-context retrieval abilities
Next, as expected, the window size is an important parameter for the performance of HQLT on these tasks overall. However, breaking down the performance on the task level, there are mixed results. For example, increasing the window size from 64 to 128 yields large degradation on SWDE (from 20.0 down to 16.9), while increasing it up to 512 yields an effective improvement (SWDE accuracy going up to 22.9). Similarly, increasing the window size from 64 to 128 yields a large improvement on FDA (from 10.9 to 17.9) but increasing it further to 256 yields degradation (down to 15.1). These observations confirm previously reported sensibility of these tasks [6]. Nevertheless, overall, we observe gradual performance increase on average by increasing the window size from the default value of 64 to 512, and finally up to 1024. However, this performance gain is not for free: increasing the window size increases the compute requirement for HQLTs (see further discussion on this compute-performance trade off in Appendix B.1). Finally, by increasing the parameter count to 1.3B, HQLT with a tiny window size of 64 still yields a large improvement over the standalone DeltaNet on FDA.
Figure 2: Evaluation of the Synchronous Hybrid Quadratic-Linear Transformer (HQLT) on reinforcement learning in partially observable environments, using a “passive visual match” task [42, 43]. A: In this task, an agent (the beige pixel) navigates in a 2D grid world (of size $7\times 11$) delimited by impermeable walls (black). The agent can only observe the nearby pixels ($5\times 5$-grid centered on the agent; illustrated by the light-blue boxes). An episode in this task has three phases. During Phase 1 (whose duration is 15 time steps), the agent observes a color, randomly drawn from three choices, red, green or blue (here blue). In Phase 2 (750 steps), the agent is in a room with apples (green); collecting an apple yields a reward of 1. There are initially 10 apples, and they reappear every 20 steps; their positions are random. In Phase 3 (max. 15 steps), if the agent reaches the pixel with the color that matches the one provided in Phase 1, the episode ends successfully; it yields a reward of 100. Alternatively, Phase 3 terminates if the agent reaches a pixel with a wrong color or when the limit of 15 steps is reached (no reward is given in these cases). B and C show the average return and success rate over 20 test episodes as a function of training environment steps, respectively. We show the average and 95% confidence intervals computed using three training seeds. The variation in success rate is high for Transformer, as one of the seeds consistently achieved the 100% success rate after certain training steps, while other seeds did not. Similarly for HQLT, one of the seeds consistently achieved above 70%.

### v2

@@ Abstract
We develop hybrid memory architectures for general-purpose sequence processing neural networks, that combine key-value memory using softmax attention (KV-memory) with fast weight memory through dynamic synaptic modulation (FW-memory)—the core principles of quadratic and linear transformers, respectively. These two memory systems have complementary but individually limited properties: KV-memory offers precise retrieval but is constrained by quadratic complexity in sequence length, while FW-memory supports arbitrarily long sequences and enables more expressive computation but sacrifices precise recall. We propose and compare three methods to blend these two systems into a single memory system, differing in how and when input information is delivered to each system, to leverage the strengths of both. We conduct experiments on general language modeling and retrieval tasks by training 340M- and 1.3B-parameter models from scratch, as well as on synthetic algorithmic tasks designed to precisely illustrate the benefits of certain hybrid methods over others. We also evaluate our hybrid memory systems on reinforcement learning in partially observable environments. Overall, we demonstrate how a well-designed hybrid can overcome the limitations of its individual components, offering new insights into the design principle of neural memory systems. [^1] Our code is public: https://github.com/kazuki-irie/hybrid-memory.
@@ 1 Introduction
Modern transformers come in two flavors. The standard or “quadratic transformer” (QT) [1] leverages softmax attention over explicit key-value memory storage which enables precise memory retrieval [2], while the length of sequences they can handle is limited in practice due to its quadratic computational complexities (as reflected in our model terminology). On the other hand, “linear transformers” (LT), also known as fast weight programmers (FWPs) [3, 4] sacrifice the precision of the softmax function (the complexity bottleneck of QT), but in return they support the processing of arbitrarily long sequences. Furthermore, some of their variants (called DeltaNet models [5, 6, 7, 8]) enable more expressive computations (e.g., certain types of state-tracking), which QTs struggle to perform [7, 8, 9, 10, 11, 12, 13, 14, 15, 16].
@@
We propose and compare three blending methods that are conceptually well-motivated—they differ in how and when input information is delivered to each memory system, and empirically evaluate them to ultimately converge on a single architecture. We conduct experiments to test general language modeling and in-context retrieval abilities (using the standard lm-evaluation-harness [19]), by training 340M- and 1.3B-parameter language models from scratch using 15B tokens of the HuggingFace FineWeb-Edu dataset [20]. We also conduct experiments on synthetic algorithmic tasks (parity and modular arithmetics) [7] to determine the type of hybrid methods that can preserve the expressivity advantages of DeltaNet (which we use as the LT/FWP component of our hybrid)—a critical discussion and evaluation that are completely absent in prior discussions on quadratic and linear transformer hybrids [21, 22]. Overall, we demonstrate the benefits of hybrid model designs that can take the full advantage of the latest LT/FWP models, and show how a well-designed hybrid can mitigate the limitations of its individual components; this offers novel insights into the design principle of neural memory systems.
Finally, all our models are scalable as they are compatible with the algorithms and implementations for highly efficient quadratic and linear transformers that have been proposed in prior work on flash-attention [23] and flash-linear-attention [24] (see our code link on page 1).
@@ 2.1 Key-value memory & Softmax attention
where ${\bm{q}}_{t},{\bm{k}}_{t}\in\mathbb{R}^{d_{\text{key}}}$, ${\bm{v}}_{t}\in\mathbb{R}^{d_{\text{out}}}$ (with $[]$ denoting vector concatenation), ${\bm{W}}^{\text{slow}}\in\mathbb{R}^{(2d_{\text{key}}+d_{\text{out}})\times d_{\text{in}}}$ is the trainable weight matrix (meaning of the superscript “slow” becomes clear in Sec. 2.2), $\oplus$ as in ${\bm{K}}_{t-1}\oplus{\bm{k}}_{t}$ denotes concatenation of vector ${\bm{k}}_{t}\in\mathbb{R}^{d_{\text{key}}}$ to matrix ${\bm{K}}_{t-1}\in\mathbb{R}^{d_{\text{key}}\times(t-1)}$ along the time dimension, yielding ${\bm{K}}_{t}\in\mathbb{R}^{d_{\text{key}}\times t}$ (${\bm{K}}_{0}$ and ${\bm{V}}_{0}$ are initially empty). We omit the $1/\sqrt{d_{\text{key}}}$ scaling inside $\mathrm{softmax}$, as well as the output projection, which are irrelevant to our discussion.
Eq. 3 is the so-called softmax attention operation; the query vector from the current step $t$ is compared to the keys from all the time steps through dot product; these similarity scores are sharpened through the softmax function (crucial for precise retrieval), and the resulting scores are used as coefficients to compute the weighted average of all the value vectors to produce the output. As can be seen in Eqs. 2, the sizes of key and value memory storage linearly grow with the time step (i.e., sequence length), resulting in quadratic complexity w.r.t. sequence length in the attention computation in Eq. 3. Consequently, practical self-attention requires predetermining a maximum context length, i.e., size of the sliding window, and any input that falls outside the window is discarded. On the other hand, training of such a model can be highly efficient on modern hardware as the computation above is parallelizable over time steps/sequence elements.
@@ 2.2 Fast weight memory & Linear attention and delta rule
where the dimension of ${\bm{W}}^{\text{slow}}$ is increased to be $\in\mathbb{R}^{(2*d_{\text{key}}+d_{\text{out}}+1)\times d_{\text{in}}}$, i.e., one extra output dimension (+1+1) is introduced to produce an addition variable $\beta_{t}\in\mathbb{R}$. $\sigma$ is an activation function applied to $\beta_{t}$, which is typically sigmoid or $2$ times sigmoid (the factor $2$ introduces negative eigenvalues in the state transition matrix enabling strong state-tracking abilities [7]). The choice of $\phi$ is particularly important for stability when the delta rule is used [5]; here we define $\phi$ as element-wise sigmoid linear unit (SiLU; $x$ times sigmoid) followed by $L_{2}$ normalization as in Yang et al. [29]. Eq. 8 corresponds to a rank-one update of the fast weight matrix, from ${\bm{W}}_{t-1}$ to ${\bm{W}}_{t}$, through the delta learning rule [34], where the slow net-generated variables, ${\bm{v}}_{t}$, $\phi({\bm{k}}_{t})$, and $\sigma(\beta_{t})$, play the role of target, input, and learning rate of the delta rule, respectively.
@@ 3.1 Delayed-Streaming HQLT
Let $\ominus$ denote the ‘remove’ operation taking a matrix and one of its column vectors to be removed as arguments. $S$ denotes the window size, i.e., the number of key-value pairs KV-memory can store.
At every time step $t$, an HQLT takes an input ${\bm{x}}_{t}\in\mathbb{R}^{d_{\text{in}}}$, and produces an output ${\bm{y}}_{t}\in\mathbb{R}^{d_{\text{out}}}$, while maintaining two types of memory: (1) KV-memory, denoted as two matrices ${\bm{K}}_{t-1}\in\mathbb{R}^{d_{\text{key}}\times S}$ and ${\bm{V}}_{t-1}\in\mathbb{R}^{d_{\text{out}}\times S}$, and (2) FW-memory, denoted as a matrix ${\bm{W}}_{t-1}\in\mathbb{R}^{d_{\text{out}}\times d_{\text{key}}}$. As in DeltaNet (Sec. 2.2), the trainable parameter matrix is ${\bm{W}}^{\text{slow}}\in\mathbb{R}^{(2*d_{\text{key}}+d_{\text{out}}+1)\times d_{\text{in}}}$, and variables are first generated from the input as follows:
@@ 3.2 Delayed-Chunk HQLT
While Delayed-Streaming HQLT above seems natural from the perspective of streaming causal sequence processing, another “delayed” variation can be motivated by the chunk-wise training algorithm of Eq. 6 (left). In fact, an HQLT can be naturally obtained by introducing the softmax activation in the intra-chunk attention operation of Eq. 6 (left; the second term), i.e., $(\mathbf{K}_{\mathbf{n}}^{\top}\mathbf{Q}_{\mathbf{n}}\odot\mathbf{M})$; and by using this chunk-wise processing for inference too. We call this approach “Delayed-Chunk HQLT”. We provide further comments about this model in Appendix B.1.
@@ 4.1 Evaluating general language modeling capabilities
We begin with evaluating general language modeling capabilities of the proposed HQLT models as a general sanity check. We train language models with either 340M or 1.3B trainable parameters (using sequence lengths of 2048 and 2240, respectively, unless otherwise noted) from scratch on 15B tokens of the HuggingFace FineWeb-Edu dataset [20]. These choices mostly follow the configurations of the recent work on LTs/DeltaNet [28, 29, 7] and their most recent recommendations available on the flame repository [24], including the choice of tokenizer (fla-hub/transformer-1.3B-100B). The subword-unit vocabulary size is 32K for all models. All models have 24 layers. The baseline Transformer architecture is from [38] (denoted as Transformer++, following prior convention) and the DeltaNet configuration is from [29]; both models/configurations are used in prior work [29, 7]. Note, however, that for fair comparison with the baseline quadratic transformer, we do not use short-window convolution in the DeltaNet, unlike in recent work [29].
@@
       window $\times 2$ = 128 26.6 27.8 35.6 68.1 41.7 51.2 61.4 30.1 48.0
       window $\times 4$ = 256 26.7 27.7 35.4 67.0 42.2 52.7 59.8 29.0 47.7
       window $\times 8$ = 512 27.0 28.0 35.9 66.3 41.3 53.2 60.1 29.0 47.6
@@ 4.2 Evaluating Expressivity
Here we compare HQLT variants’ abilities to leverage expressivity advantages of the FW-memory component, DeltaNet. We evaluate models on two popular regular language recognition tasks, parity and modular arithmetic without brackets, which are well-known tasks that quadratic transformers and other linear recurrent models fail (due to lack of expressivity), while more expressive, DeltaNet performs well [7]. We mainly follow Grazzi et al. [7]’s experimental settings: we use 2-layer and 3-layer models for parity and modular arithmetic, respectively; training and test sequences have lengths from 3 up to 40, and from 40 to 256, respectively. The window size for all HQLTs is set to 8 (except for Delayed-Chunk which uses 16). We refer to Appendix A.2 for further experimental details.
Table 3 shows the results. As expected, the two “delayed” variants, Delayed-Stream and Delayed-Chunk, fail on these tasks just like the quadratic transformer. In these “delayed” variations, FW-memory has a delay of a few time steps to access an input, corresponding to the quadratic attention window size; this could be too late to leverage FW-memory’s capabilities for these tasks requiring state-tracking in “real time”. In contrast, Synchronous HQLT has no such limitation and successfully solves these tasks.
@@
Given that the delayed versions have no other significant advantage over the synchronous version (apart from the conceptual elegance in introducing explicit division of labors over time), we consider this result as a strong argument to prefer the Synchronous variant over other blending strategies for HQLT. This also allows us to conclude that the design of “Infini-attention” from prior work [22], which makes use of the Delayed-Chunk strategy, is sub-optimal in light of expressivity.
Table 4: Evaluating Hybrid Quadratic-Linear Transformers (HQLTs) on retrieval-intensive tasks. Both the 340M/1.3B models are trained for 15B tokens. The default window size for KV-memory is 64 in HQLTs. Dynamic mixers (scalar and vector) add 0.4M and 25M parameters, respectively.
Window SWDE SQuAD FDA Avg.
size acc $\uparrow$ acc $\uparrow$ acc $\uparrow$
@@
     sum mixer 64 13.3 26.2 12.6 17.4
     dynamic scalar 64 21.1 27.7 11.4 20.1
     dynamic vector 64 20.0 28.4 10.9 19.8
       window $\times 2$ 128 16.9 30.5 17.9 21.8
       window $\times 4$ 256 18.6 35.6 15.1 23.1
       window $\times 8$ 512 22.9 35.7 17.3 25.3
       window $\times 16$ 1024 34.0 37.1 50.1 40.4
@@
Transformer++ 2048 53.7 41.5 64.7 53.3
DeltaNet - 32.9 29.9 23.6 28.8
HQLT Synchronous 64 31.9 31.2 30.2 31.1
@@ 4.3 Evaluating in-context retrieval abilities
Next, as expected, the window size is an important parameter for the performance of HQLT on these tasks overall. However, breaking down the performance on the task level, there are mixed results. For example, increasing the window size from 64 to 128 yields large degradation on SWDE (from 20.0 down to 16.9), while increasing it up to 512 yields an effective improvement (SWDE accuracy going up to 22.9). Similarly, increasing the window size from 64 to 128 yields a large improvement on FDA (from 10.9 to 17.9) but increasing it further to 256 yields degradation (down to 15.1). These observations confirm previously reported sensibility of these tasks [6]. Nevertheless, overall, we observe gradual performance increase on average by increasing the window size from the default value of 64 to 512, and finally up to 1024. Another interesting comparison is between HQLT with a window size of 1024 and the baseline transformer trained with the same window size (second row); we obtain a large improvement (40.4 vs. 29.0) showing the benefit of an extra FW-memory that covers a larger context, even though retrieval is not the main strength of FW-memory. Finally, by increasing the parameter count to 1.3B, HQLT with a tiny window size of 64 still yields a large improvement over the standalone DeltaNet on FDA. The compute requirements for HQLTs are discussed in Appendix A.1 and B.2.
Figure 2: Evaluation of the Synchronous Hybrid Quadratic-Linear Transformer (HQLT) on reinforcement learning in partially observable environments, using a “passive visual match” task [48, 49]. A: In this task, an agent (the beige pixel) navigates in a 2D grid world (of size $7\times 11$) delimited by impermeable walls (black). The agent can only observe the nearby pixels ($5\times 5$-grid centered on the agent; illustrated by the light-blue boxes). An episode in this task has three phases. During Phase 1 (whose duration is 15 time steps), the agent observes a color, randomly drawn from three choices, red, green or blue (here blue). In Phase 2 (750 steps), the agent is in a room with apples (green); collecting an apple yields a reward of 1. There are initially 10 apples, and they reappear every 20 steps; their positions are random. In Phase 3 (max. 15 steps), if the agent reaches the pixel with the color that matches the one provided in Phase 1, the episode ends successfully; it yields a reward of 100. Alternatively, Phase 3 terminates if the agent reaches a pixel with the wrong color or when the limit of 15 steps is reached (no reward is given in these cases). B and C show the average return and success rate over 20 test episodes as a function of training environment steps, respectively. We show the average and 95% confidence intervals computed using three training seeds. The variation in success rate is high for Transformer, as one of the seeds consistently achieved the 100% success rate after certain training steps, while other seeds did not. Similarly for HQLT, one of the seeds consistently achieved above 70%.